OpenStudy (anonymous):

plz i want to know step by step how to do this. i don't want direct answers, thanks i know the answer too, just wanna review this Solve for x : 2/3 (x - 4) = 2x

OpenStudy (trojanpoem):

at first try to get rid of the fraction by hitting both sides with 3/2 now we will have x-4 = 3x now simply take all x's to side and numbers to side 2x = -4 x = -4 /2 x = -2

OpenStudy (anonymous):

hitting both sides by 3/2 means multiplying it by that? kinda need t know lol

OpenStudy (anonymous):

cause 2 x 3/2 = 3 so do you do that anytime when you have a problem like this

OpenStudy (trojanpoem):

Yep , I love using strange math expressions

OpenStudy (anonymous):

OK i did it a different way and still got -2 thanks

OpenStudy (trojanpoem):

Something like this ? 2x/ 3 - 8/3 = 2x 4x/3 = -8/3 x =2 ?

OpenStudy (anonymous):

yes

OpenStudy (trojanpoem):

Totally right.